OpenStudy (anonymous):

Evaluate and write your solution in scientific notation:(4.1 x 105)(2 x 107) A. 8.2 x 1012 B. 8.2 x 10-2 C. 8.2 x 1035

jimthompson5910 (jim_thompson5910):

Multiply the base numbers 4.1 and 2 to get...?

OpenStudy (anonymous):

8.2 :D

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

now add the exponents: 5 and 7 to get ???

OpenStudy (anonymous):

35?

jimthompson5910 (jim_thompson5910):

no you're adding the exponents, not multiplying them

OpenStudy (anonymous):

*12 sorry xD misread somehow

jimthompson5910 (jim_thompson5910):

think of it like this x^2 times x^3 is x^(2+3) = x^5

jimthompson5910 (jim_thompson5910):

You got it. So the answer is \[\Large 8.2 \times 10^{12}\]

OpenStudy (anonymous):

Thanks :D

jimthompson5910 (jim_thompson5910):

yw